# Stealing That Free Lunch: Exposing the Limits of Dyna-Style Reinforcement Learning

### v2

@@ 4.2 What If We Had a Perfect Model?
In OpenAI Gym, even MBPO with a learned model achieves sample efficiency comparable to or exceeding SAC across all environments (cf. Figure 3). However, in DMC, while a perfect model enables policy improvement in environments where MBPO with a learned ensemble fails completely, it does not consistently surpass SAC’s sample efficiency in three out of the six environments shown in Figure 9.
@@ 4.3 Investigating Underlying Failure Mechanisms Induced By Model Bias
Overestimation, common in offline and online off-policy methods with high replay ratios, arises when a learned Q-function is queried for state-action pairs that are out-of-distribution for the training data (Thrun & Schwartz, 1993). Online methods typically mitigate this by using real experiences to exploit high-reward regions, naturally correcting overestimation through interaction. Conversely, underestimation occurs in unseen regions where the Q-function is overly pessimistic. Unlike overestimation, underestimation is harder to address, as it requires purely exploratory actions to visit these regions, which policies inherently avoid due to their focus on maximizing expected return.
@@
To validate this hypothesis, we compared the Q-values obtained using both a learned model and a perfect model throughout training. [^4] These results are in Appendix F in Figure 16. With a perfect model, massive underestimation essentially vanished, implicating modeling errors as the primary driver of this issue. However, per the discussion of Section 4.2, even with a perfect model and no underestimation MBPO fails to match SAC’s performance.
Since critic divergence stems from learned functions extrapolating in an unconstrained manner (Ball et al., 2023), we employed Layer Normalization (Ba et al., 2016a) as a regularization technique as in (Ball et al., 2023) to bound the critic output. This technique has proven quite successful in previous work (Nauman et al., 2024), but did not mitigate the critic divergence induced by the non-stationarity of the critic target as shown in Appendix F in Figure 17.
Key Insights: MBPO’s learned model struggles to make accurate predictions in DMC, even on the training distribution. A perfect 1-step model improves MBPO in DMC but still fails to match or outperform SAC, which was previously achieved in Gym. Predictive model errors and non-stationary critic targets induce critic divergence. Layer Normalization fails to mitigate critic instability caused by model errors in MBPO.
@@ 5 If It’s Not Just The Model Could It Be The Learning Dynamics?
Our analysis in Section 4 suggests that errors in the predictive model contribute to MBPO’s poor performance in DMC, but cannot fully explain the issues identified in Section 3. This observation prompts us to explore other potential contributing factors, and in this section we investigate the interplay between learning dynamics and the plasticity of the networks that make up MBPO, seeking to understand if these factors might hinder either the predictive model’s accuracy or the base off-policy method’s ability to improve policy performance.
@@
Plasticity loss — where prolonged training diminishes a network’s capacity to learn new tasks — has been studied in off-policy RL (Lyle et al., 2022) and model-based RL (Qiao et al., 2024), with successful applications in DMC (Nikishin et al., 2022; D’Oro et al., 2023). High replay ratios in off-policy RL are known to exacerbate plasticity issues, as both the predictive model and the learned Q-function face continually changing data distributions and using a high replay ratio forces them to learn to solve a sequence of similar, but distinct, tasks (Dabney et al., 2021).
@@
Qiao et al. (2024) demonstrate that periodic reinitialization of the learned model parameters can mitigate the loss of plasticity in model-based RL and enhance model accuracy. Therefore, to determine if a loss of plasticity is contributing to MBPO’s failures, we completely reset all parameters of the predictive model every $2\times 10^{4}$ environment steps; however, we found that this procedure yields identical performance to that of Figure 4.555Training curves may be found in Figure 18 in Appendix F. These results indicate that plasticity loss in the predictive model is not a contributing factor to the low model accuracy of MBPO.
@@
Next, we investigated whether plasticity issues were affecting the networks that comprise the off-policy base of MBPO — the actor, critic, target critic, and the automatically tuned temperature in SAC. As for the investigation of the model loss of plasticity, we reset all of the aforementioned networks, except the predictive model, every $2\times 10^{4}$ interactions. In Figure 11, we see that in both quadruped tasks performing parameter resets of the off-policy base neural networks allows MBPO to not only improve its policy from the initial random policy, but outperform SAC by a large margin. Additionally we see that in the humanoid-walk task performing periodic resets allows MBPO to improve from the initial random policy, but it still underperforms SAC. In all other tasks MBPO with periodic resets still cannot improve its policy beyond the initial random policy. We can conclude that plasticity issues related to these base off-policy components do, at least partially, account for some failings of MBPO.
@@
Figure 11: Comparing performance of MBPO and SAC with and without periodic resets of the the actor, critic, target critic, and the automatically tuned temperature for the six tasks. Adding periodic resets to MBPO allows it to improve the policy performance beyond that of the random initial policy for some environments, but when compared to SAC with resets MBPO performs much worse across all tasks. Full results are in Appendix F in Figure 13.
@@
Nevertheless, alleviating plasticity loss is not a universal solution. Of even more significance, the resets we apply for MBPO are equally applicable for SAC (i.e., MBPO without Dyna-based “enhancements”), and when applying periodic resets to SAC we see in Figure 11 that SAC significantly outperforms MBPO with periodic resets across most tasks, which further shows that Dyna-style enhancements actually degrade the performance of their off-policy base algorithm in DMC.
@@ 6 Conclusion
Recognizing that the techniques we applied are not exhaustive, we also aim to foster collaboration and transparency by including a frequently asked questions discussion in Appendix A. This section addresses common questions and provides additional context for our work. We intend to update it continually, incorporating insights and feedback from the broader research community as we refine our methods.
7 Acknowledgements
We would like to thank Harshit Sikchi, Eugene Vinitsky, Siddhant Agarwal, and Max Rudolph for insightful discussions and their feedback on this paper.
@@ Appendix A Frequently Asked Questions
No. As mentioned at the end of Section 2.1, DreamerV3 (Hafner et al., 2024) achieves state-of-the-art performance in DMC environments. This evidence suggests that there are two subclasses of DMBRL algorithms: those that succeed in DMC and those that do not. This paper highlights two algorithms in the latter subclass, with the aim of inspiring future work to identify algorithmic improvements that enable all DMBRL algorithms to perform successfully not only in OpenAI Gym or DMC, but across a broader range of tasks.
@@
Yes. We did not find any significant overfitting of the model in either the OpenAI Gym environments or in DMC. Further, we performed intermittent model resets in Section 5 as a means to mitigate this issue if overfitting was occurring throughout training (Qiao et al., 2024). As discussed in Section 5, resets do not resolve the performance gap.
@@
This paper does not (and cannot) exhaustively consider all potential of sources of the performance gap we observed. If you can think of additional experiments to try, we would love to hear from you via the email provided at the bottom of the first page.
@@ Appendix B Timing Claim for MBPO Compared to Our Implementation
Based on previous work (Ghugare et al., 2023), MBPO’s average time per environment step was approximately $0.6$ seconds across Gym environments. Even if we optimistically reduce this to $0.5$ seconds per step (to account for our usage of slightly more modern computational hardware) the time required to run $6$ seeds across the $6$ DMC environments shown in Figure 4 for $500$k environment steps is: $\frac{0.5\times 500000\times 6\times 6}{(3600\times 24)}$ $=$ $104$ days.
@@ Appendix C Extended Discussion on Benchmark Modeling Differences
As discussed in Section 2.2, OpenAI Gym and DM Control (DMC) use different reward structures, physical parameters, and termination conditions. These differences certainly could contribute to Dyna-style algorithms failing in DMC but not Gym. However, both are equally conventional and widely accepted testbeds for reinforcement learning. This is evidenced by the fact that DMC has been successfully used for demonstrating RL with not only proprioceptive observations (D’Oro et al., 2023; Nikishin et al., 2022), but for visual observations too (Laskin et al., 2020; Hafner et al., 2020; Srinivas et al., 2020; Yarats et al., 2021b).
It is important to emphasize that our goal in this work is not to identify what specific aspects of the environment might cause DMBRL to fail. Instead, we aim to show that DMBRL performs reasonably well on a fair and rigorous benchmark, underscoring its robustness and applicability.
@@
Conversely, MBPO performed relatively well on walker tasks in both Gym and DMC environments, as shown in Figures 3 and 9, despite DMC lacking termination conditions. Interestingly, ALM exhibited a different pattern, failing on walker in DMC but succeeding in Gym. Additionally, switching frameworks reversed the relative performance of SAC and MBPO for walker. In the humanoid environment, both MBPO and ALM succeeded in Gym but failed entirely in DMC. These results indicate that termination conditions alone do not account for DMBRL method failures across multiple environments that are broadly similar, as performance varies across frameworks and tasks.
@@ D.1 ALM
The code to run ALM was used with only minimal modifications introduced to allow interfacing with DMC environments. No hyperparameters were changed in either our Gym experiments or in our DMC experiments from what was provided in the original ALM repository. When applying ALM without Dyna-style enhancements we simply replaced the lambda svg loss with the td loss that is provided in the experimental branch of the ALM repository and kept all hyperparameters fixed.
We implement ALM using DDPG (Lillicrap et al., 2019) as the base algorithm. Following prior SVG methods (Amos et al., 2020), we parameterize the encoder, model, policy, reward, classifier and Q-function as 2-layer neural networks, all with 512 hidden units except the model which has 1024 hidden units. The model and the encoder output a multivariate Gaussian distribution over the latent-space with diagonal covariance. Like prior work (Hansen et al., 2022; Yarats et al., 2021a), we apply layer normalization (Ba et al., 2016b) to the value function and rewards. Similar to prior work (Schulman et al., 2015; Hafner et al., 2019), we reduce variance of the policy objective by computing an exponentially-weighted average of the objective for rollouts of length 1 to an integer $K$ chosen as a hyperparameter. To train the policy, reward, classifier and Q-function we use the representation sampled from the target encoder. For exploration, we use added normal noise with a linear schedule for the standard deviation (Yarats et al., 2021a). All hyperparameters are listed in Table 3. The specific implementation details below are reproduced with minor modification from (Ghugare et al., 2023):
For additional details see (Ghugare et al., 2023).
@@ D.2 MBPO
Code is forthcoming.
Appendix E Hyperparameters for SAC, MBPO, and ALM
@@
Table 3: Hyperparameters used for ALM with and without Dyna-style enhancements and reproduced from (Ghugare et al., 2023).
@@
Appendix F Full Experimental Results

### v3

@@ 4.2 What If We Had a Perfect Model?
In OpenAI Gym, even MBPO with a learned model achieves sample efficiency comparable to or exceeding SAC across all environments (cf. Figure 3). However, in DMC, while a perfect model enables policy improvement in environments where MBPO with a learned ensemble fails completely, it does not consistently surpass SAC’s sample efficiency in four out of the six environments shown in Figure 6.
@@ 4.3 Investigating Underlying Failure Mechanisms Induced By Model Bias
Overestimation, common in offline and online off-policy methods with high replay ratios, arises when a learned Q-function is queried for state-action pairs that are out-of-distribution for the training data (Thrun & Schwartz, 1993). Online methods typically mitigate this by using real experiences to explore high-reward regions, naturally correcting overestimation through interaction. Conversely, underestimation occurs in unseen regions where the Q-function is overly pessimistic. Unlike overestimation, underestimation is harder to address, as it requires purely exploratory actions to visit these regions, which policies inherently avoid due to their focus on maximizing expected return.
@@
To validate this hypothesis, we compared the Q-values obtained using both a learned model and a perfect model throughout training. [^4] These results are in Appendix G in Figure 18. With a perfect model, massive underestimation is much less pronounced, implicating modeling errors as the primary driver of this issue. However, per the discussion of Section 4.2, even without massive underestimation MBPO fails to match SAC’s performance.
Since critic divergence stems from learned functions extrapolating in an unconstrained manner, we employed Layer Normalization (Ba et al., 2016a) as a regularization technique as in (Ball et al., 2023) to bound the critic output. This technique has proven quite successful in previous work (Nauman et al., 2024), but even after partly mitigating critic divergence, MBPO still cannot outperform SAC consistently as shown in Appendix G in Figure 20.
Key Insights: MBPO’s learned model struggles to make accurate predictions in DMC, even on the training distribution. A perfect 1-step model improves MBPO in DMC but still fails to match or outperform SAC, as was previously achieved in Gym. Predictive model errors and non-stationary critic targets induce critic divergence. Layer Normalization mitigates critic divergence, but does not allow MBPO to succeed.
@@ 5 If It’s Not Just The Model Could It Be The Learning Dynamics?
Section 4 suggests that errors in the predictive model contribute to MBPO’s poor performance in DMC, but cannot fully explain the issues identified in Section 3. This observation leads us to investigate how learning dynamics and network plasticity in MBPO might impact the predictive model or the base off-policy method’s abilities.
@@
Plasticity loss — where prolonged training diminishes a network’s capacity to learn new tasks — has been studied in off-policy RL (Lyle et al., 2022) and model-based RL (Qiao et al., 2023), with successful applications in DMC (Nikishin et al., 2022; D’Oro et al., 2023). High replay ratios in off-policy RL are known to exacerbate plasticity issues, as both the predictive model and the learned Q-function face continually changing data distributions and using a high replay ratio forces them to learn to solve a sequence of similar, but distinct, tasks (Dabney et al., 2021).
@@
Qiao et al. (2023) demonstrate that periodic reinitialization of the learned model parameters can mitigate the loss of plasticity in model-based RL and enhance model accuracy. Therefore, to determine if a loss of plasticity is contributing to MBPO’s failures, we completely reset all parameters of the predictive model every $2\times 10^{4}$ environment steps, which is a frequency aligned with recommendations from previous work (Qiao et al., 2023; D’Oro et al., 2023) for a replay ratio of 128. We also experimented with both more conservative intervals (e.g., every $1.28\times 10^{5}$ environment steps) and intermediate values, but found that results were indistinguishable from those in Figure 4 regardless of reset interval. These findings suggest that model plasticity is not the primary cause of low model accuracy in MBPO. [^5] Training curves for model resets every $2\times 10^{4}$ environment steps may be found in Figure 20 in Appendix G.
@@
Next, we investigated whether plasticity issues were affecting the networks that comprise the off-policy base of MBPO — the actor, critic, target critic, and the automatically tuned temperature in SAC. Unlike the investigation of the model losses due to plasticity, we reset all of the aforementioned networks, except the predictive model, every $2\times 10^{4}$ interactions. In Figure 8, we see that in both quadruped tasks performing periodic resets allows MBPO to not only improve its policy from the initial random policy, but outperform SAC by a large margin. Additionally we see that in the humanoid-walk task performing periodic resets allows MBPO to improve from the initial random policy, but it still underperforms SAC. In all other tasks MBPO with periodic resets still cannot improve its policy beyond the initial random policy. We can conclude that plasticity issues related to these base off-policy components do, at least partially, account for some failings of MBPO.
@@
Figure 8: Comparison of MBPO and SAC with and without periodic resets of the actor, critic, target critic, and automatically tuned temperature. While periodic resets help MBPO improve policy performance beyond the random initial policy in some environments, SAC with resets consistently outperforms MBPO across all tasks. Full results are in Appendix G and Figure 15.
@@
Nevertheless, alleviating plasticity loss is not a universal solution. Of even more significance, the periodic resets applied in MBPO are equally applicable to SAC. As shown in Figure 8, SAC with periodic resets significantly outperforms MBPO with the same resets on most tasks. This highlights that Dyna-style enhancements can degrade the performance of their off-policy base algorithm in DMC.
@@ 6 Conclusion
Recognizing that our work is not exhaustive, we also have included a frequently asked questions section in Appendix A to address common questions, provide context, and foster collaboration. We intend to update it continually, incorporating insights and feedback from the broader research community as we refine our methods.
Acknowledgements
This work was supported by the National Science Foundation under Grant No. 2409535. We would like to thank Harshit Sikchi, Eugene Vinitsky, Siddhant Agarwal, and Max Rudolph for insightful discussions and their feedback on this paper.
Impact Statement
This paper underscores the importance of robust and diverse benchmarking in the Machine Learning community to ensure that reported progress translates to meaningful advances. By pointing to the limitations of Dyna-style methods in broader contexts, we encourage the community to critically re-evaluate widely adopted approaches and the systems that reward them. Addressing systemic issues, such as the incentives for replicating influential results without deeper scrutiny, is essential to fostering meaningful progress in the field.
Furthermore, our work highlights the barriers to accessibility in reproducibility and evaluation. Many influential methods, including those analyzed in this study, remain difficult to reproduce due to the significant computational resources required to replicate results or run new experiments. Such challenges prevent broader participation in the refinement and critical evaluation of algorithms, especially from researchers with limited access to cutting-edge hardware. To help address these concerns for Dyna-style methods, we provide efficient, open-source code that significantly reduces evaluation costs and facilitates more inclusive collaboration within the research community.
From an ethical standpoint, this work does not present immediate societal risks, but rather focuses on improving the methodological rigor within reinforcement learning research. By advocating for more comprehensive evaluation practices, reducing resource requirements, and fostering transparency, we hope to encourage accountability and accessibility in research practices.
We believe this work holds value for the community by drawing attention to challenges in generalization, overfitting to benchmarks, accessibility, and the broader need for a culture that prioritizes the evaluation of actual progress. While our study is not exhaustive, we provide a framework for further investigation and invite collaboration to refine and build upon our findings. By enabling the community to address these shortcomings, we aim to facilitate research that leads to more robust, impactful, and accessible algorithms.
@@ Appendix A Frequently Asked Questions
No. As mentioned at the end of Section 2.1, DreamerV3 (Hafner et al., 2025) achieves state-of-the-art performance in DMC environments. This evidence suggests that there are two subclasses of DMBRL algorithms: those that succeed in DMC and those that do not. This paper highlights two algorithms in the latter subclass, with the aim of inspiring future work to identify algorithmic improvements that enable all DMBRL algorithms to perform successfully not only in OpenAI Gym or DMC, but across a broader range of tasks.
@@
Yes. We did not find any significant overfitting of the model in either the OpenAI Gym environments or in DMC. Further, we performed intermittent model resets in Section 5 as a means to mitigate this issue if overfitting was occurring throughout training (Qiao et al., 2023). As discussed in Section 5, resets do not resolve the performance gap.
@@
This paper does not (and cannot) exhaustively consider all potential of sources of the performance gap we observed. If you can think of additional experiments to try, we would love to hear from you.
@@ Appendix B Timing Claim for MBPO Compared to Our Implementation
Based on previous work (Ghugare et al., 2022), MBPO’s average time per environment step was approximately $0.6$ seconds across Gym environments. Even if we optimistically reduce this to $0.5$ seconds per step (to account for our usage of slightly more modern computational hardware) the time required to run $6$ seeds across the $6$ DMC environments shown in Figure 4 for $500$k environment steps is: $\frac{0.5\times 500000\times 6\times 6}{(3600\times 24)}$ $=$ $104$ days.
@@ Appendix C Extended Discussion on Benchmark Modeling Differences
As discussed in Section 2.2, OpenAI Gym and DM Control (DMC) use different reward structures, physical parameters, and termination conditions. These differences certainly could contribute to Dyna-style algorithms failing in DMC but not Gym. However, both are equally conventional and widely accepted testbeds for reinforcement learning. This is evidenced by the fact that DMC has been successfully used for demonstrating RL with not only proprioceptive observations (D’Oro et al., 2023; Nikishin et al., 2022), but for visual observations too (Laskin et al., 2020; Hafner et al., 2020; Laskin et al., 2020; Yarats et al., 2022).
It is important to emphasize that the goal of this work is not to pinpoint specific environmental factors that might cause DMBRL to fail. Instead, we aim to demonstrate that while DMBRL performs well on one benchmark, it largely fails on another that, at first glance, appears quite similar. This observation raises concerns about the robustness of DMBRL algorithms. Rather than delving into the algorithmic causes of this issue as we have for the majority of this paper, this section highlights and discusses key differences between the two benchmarks, providing a foundation for future work to investigate environment-specific failures of DMBRL algorithms.
@@
Conversely, MBPO performed relatively well on walker tasks in both Gym and DMC environments, as shown in Figures 3 and 6, despite DMC lacking termination conditions. Interestingly, ALM exhibited a different pattern, failing on walker in DMC but succeeding in Gym. Additionally, switching frameworks reversed the relative performance of SAC and MBPO for walker. In the humanoid environment, both MBPO and ALM succeeded in Gym but failed entirely in DMC. These results indicate that termination conditions alone do not account for DMBRL method failures across multiple environments that are broadly similar, as performance varies across frameworks and tasks.
@@ D.1 ALM
The code to run ALM was used with only minimal modifications introduced to allow interfacing with DMC environments. No hyperparameters were changed in either our Gym experiments or in our DMC experiments from what was provided in the original ALM repository. When applying ALM without Dyna-style enhancements we simply replaced the lambda svg loss with the td loss that is provided in the experimental branch of the ALM repository and kept all hyperparameters fixed. The specific implementation details below are reproduced with minor modification from (Ghugare et al., 2022):
We implement ALM using DDPG (Lillicrap et al., 2019) as the base algorithm. Following prior SVG methods (Amos et al., 2021), we parameterize the encoder, model, policy, reward, classifier and Q-function as 2-layer neural networks, all with 512 hidden units except the model which has 1024 hidden units. The model and the encoder output a multivariate Gaussian distribution over the latent-space with diagonal covariance. Like prior work (Hansen et al., 2022; Yarats et al., 2022), we apply layer normalization (Ba et al., 2016a) to the value function and rewards. Similar to prior work (Schulman et al., 2015; Hafner et al., 2020), we reduce variance of the policy objective by computing an exponentially-weighted average of the objective for rollouts of length 1 to an integer $K$ chosen as a hyperparameter. To train the policy, reward, classifier and Q-function we use the representation sampled from the target encoder. For exploration, we use added normal noise with a linear schedule for the standard deviation (Yarats et al., 2022). All hyperparameters are listed in Table 3.
For additional details see (Ghugare et al., 2022).
@@ D.2 MBPO
The code for our MBPO implementation is available at https://github.com/CLeARoboticsLab/STFL.
Appendix E Nominal Hyperparameters for SAC, MBPO, and ALM
@@
Table 3: Hyperparameters used for ALM with and without Dyna-style enhancements and reproduced from (Ghugare et al., 2022).
@@
Appendix F Hyperparameter Sweeps
We conducted a large-scale hyperparameter sweep to evaluate MBPO’s sensitivity in hopper-stand. This sweep comprised $93$ hyperparameter configurations, with each configuration evaluated across three seeds to capture variation due to initialization and training dynamics, resulting in a total of 279 runs. The full range of values is listed in Table 4. All runs used a synthetic-to-real data ratio of $0.95$ and were trained online for $50$k environment steps. No combination of hyperparameters in this sweep yielded consistent performance.
To test whether the findings central to our claims generalized across environments and longer training durations, we conducted an additional sweep using single-seed runs on both hopper-stand and humanoid-stand. For each environment, we evaluated $81$ hyperparameter configurations for $200$k environment steps (see Table 5). These experiments targeted learning dynamics across a higher number of environment interactions and helped ensure that the lack of improvement was not simply due to undertraining or insufficient data. No combination of hyperparameters yielded consistent performance.
Since none of the $174$ hyperparameter variations across $2$ environments and $441$ runs led to meaningful performance gains, we omit per-parameter plots for brevity. Instead, Figure 9 and Figure 10 show unlabeled return curves for each configuration. Despite substantial hyperparameter sweeps, return performance remained poor, reinforcing our claim that model-based policy optimization fails to benefit from tuning in this setting.
Table 4: Hyperparameter sweep ranges for MBPO on hopper-stand.
Hyperparameter Values
ensemble_hidden {50, 100, 200, 400}
ensemble_lr {1.5e-4, 3e-4, 6e-4}
ensemble_interval {125, 250, 500}
model_train_steps {1250, 2500, 5000, 10000}
model_rollouts_per_step {100, 200, 400, 800, 1600}
num_elites {1, 3, 5}
Table 5: Hyperparameter sweep ranges used in single-seed runs on both hopper-stand and humanoid-stand.
Hyperparameter Values
ensemble_hidden {200, 400, 800}
ensemble_lr {1.5e-4, 3e-4, 6e-4}
ensemble_interval {125, 250, 500}
model_rollouts_per_step {200, 400, 800}
Figure 9: Hyperparameter sweep for MBPO trained on the hopper-stand environment across $93$ hyperparameter configurations (3 trials each, $\pm$ std). As none of these configurations yielded meaningful or consistent performance, we omit per-parameter plots for brevity. Despite one seed of one configuration temporarily reaching a return of 93, the policy immediately collapses to near-random performance after further training. For sake of comparison, after $50$k environment steps a SAC policy in this same task has an average return of $300$.
Figure 10: Hyperparameter sweep for MBPO trained on the hopper-stand and humanoid-stand environment across $81$ hyperparameter configurations each (1 trials each).
Appendix G Full Experimental Results